# The Impact of Spectroscopic Redshift Errors on Cosmological Measurements

### v1

@@ Abstract
Spectroscopic redshift errors, including redshift uncertainty and catastrophic failures, can bias cosmological measurements from galaxy redshift surveys at sub-percent level. We investigate their impact on full-shape clustering analysis using contaminated mock catalogs. We find that redshift uncertainty introduces a damping effect on the power spectrum. This damping is scale-dependent and absorbed by counterterms in the clustering model, keeping parameter biases below $5\%$ for the DESI survey. Catastrophic failures reduce the power spectrum amplitude by an approximately constant factor scaling with contamination rate $f_{c}$. While this effect is negligible for the DESI ELG populations ($f_{c}=1\%$), the slitless-like errors, combining redshift uncertainty with $f_{c}=5\%$ catastrophics, introduce significant biases in cosmological constraints. For this case, shifts from $6\%$ to $16\%$ ($\sim 2.2\sigma$ level) arise in estimating the fractional growth rate $df\equiv f/f^{\rm{fid}}$ and the log primordial amplitude $\ln(10^{10}A_{s})$. Applying a correction factor $(1-f_{c})^{2}$ on the galaxy power spectrum mitigates the bias but weakens the parameter constraints due to new degeneracies. Alternatively, fixing $f_{c}$ to its expected value during fitting successfully restores the unbiased posterior without loss in constraint. Our results indicate that for space-based slitless surveys such as Euclid, an accurate estimation of $f_{c}$ and its incorporation into the clustering model are essential to get unbiased cosmological constraints. Extending to evolving dark energy and massive neutrino cosmologies, we find that redshift errors do not bias the dark energy properties parametrized by $w_{0}$ and $w_{a}$, but lead to slightly weaker constraints on $\sum m_{\nu}$.
@@ 1 Introduction
The large-scale structure (LSS) observed by galaxy redshift surveys has played an important role in advancing our understanding of fundamental physics. Clustering features, such as baryon acoustic oscillations (BAO; [1, 2, 3, 4, 5]), and the full-shape power spectrum [6, 7, 8, 9], have been used to constrain some key quantities in cosmology, including dark energy equation of state $w$, the matter density $\Omega_{\mathrm{m}}$, the structure growth rate $f\sigma_{8}$, neutrino mass $\sum m_{\nu}$ and possible deviation from general relativity. Current leading experiments, the Dark Energy Spectroscopic Instrument [^1] DESI, desi.lbl.gov (DESI; [10, 11]), and the ongoing space telescope Euclid [^2] Euclid, cosmos.esa.int/web/euclid [12, 13] are mapping the 3-dimensional Universe by measuring redshifts for tens to hundreds of millions of galaxies. These surveys will cover over more than 15,000 $\rm deg^{2}$ of the sky, and trace the LSS back to redshift $\sim 3.5$. This large probing volume allows the clustering signal to be measured with minimal cosmic variance, leading to sub-percent precision in cosmological measurements. Given this level of precision, detailed studies on systematic errors are critical to ensure the robustness and reliability of cosmological constraints.
@@
Among various observational systematics, we focus on the impact of spectroscopic redshift errors. These errors arise from inaccuracies in the spectroscopic pipelines [14, 15, 16, 17]. They persist in data with secure redshifts, where calibration and algorithmic processing have already been applied. There are two primary types of errors: redshift uncertainty and redshift catastrophic failures. Redshift uncertainty refers to minor systematic offsets ($\Delta z\sim 10^{-4}$) in all redshift measurements, primarily caused by instrumental noise and the intrinsic width of spectral features. In contrast, redshift catastrophic failures result from the misidentification of spectral lines, leading to severely incorrect redshifts ($\Delta z\sim 10^{-2}$) in a small fraction of data. These systematic errors are essential because they can distort the observed clustering signal and potentially bias the cosmological inference. While previous studies have examined their individual impacts [18, 19, 20, 21, 22, 23], a comprehensive analysis that considers all these effects remains unexplored. This work aims to bridge that gap and explore their combined influence on cosmological inference from full-shape clustering analyses.
@@
The effect of spectroscopic redshift errors can be quantified using repeated redshift observations and incorporated into realistic galaxy mocks (e.g., [24, 25, 26, 21]). By comparing analyses on clean mocks with those contaminated by redshift errors, we can assess their impact on clustering statistics and cosmological constraints. Improved redshift calibration and increased repeat observation samples in DESI [27, 21] provide a better way to model the redshift errors and more precise studies of their impact on cosmological analyses. Additionally, space-based surveys such as Euclid use slitless spectroscopy at near-infrared to measure redshifts [28, 29]. This method requires no pre-selection of targets and is particularly effective for probing high-redshift galaxies. However, it is also more susceptible to redshift errors due to the lower spectrum resolution and the potential overlap of different spectra. These challenges have been addressed in the context of redshift interlopers for Euclid [23]. In this work, we use a hypothetical combined redshift errors to represent the conditions expected in slitless surveys. This approach provides a realistic description of the related systematics and offers practical guidance for interpreting and mitigating their impact in cosmological analyses for slitless survey.
@@
The outline of the paper is as follows: Section˜2 introduces two spectroscopic surveys, DESI and Euclid, and the spectroscopic redshift errors. Section˜3 describes how we construct our contaminated mock catalogs. Section˜4 presents the methodologies used to derive cosmological inferences from the mock data. In Section˜5, we present the impact of redshift errors on clustering statistics and how these propagate into the cosmological parameter constraints, including extensions $w_{0}w_{a}\rm{CDM}$ and $\nu\Lambda\rm{CDM}$ models. Finally, we summarize our findings and conclude in Section˜6.
@@ 2 Spectroscopic Survey and Redshift Errors
Spectroscopic surveys measure the redshifts of objects with high precision by analyzing their spectral features, such as emission and absorption lines. Two massive ongoing surveys, DESI and Euclid, are using spectroscopy to measure galaxy clustering with unprecedented accuracy. In its first three years, the ground-based DESI survey has observed over 14 million galaxies [5] and plans to extend its initial plan to 50M galaxies across $18{,}000\,\rm{deg}^{2}$. Its dark matter tracer includes Bright Galaxies (BGS) at $0.1<z<0.4$ [30], Luminous Red Galaxies (LRG) at $0.4<z<1.1$ [31], Emission Line Galaxies (ELG, mainly [O ii] emitters) at $0.8<z<1.6$ [32], and Quasi-Stellar Objects (QSO) at $0.8<z<3.5$ [33]. The Euclid mission conducts a slitless spectroscopic survey from space, aiming to observe tens of millions of galaxies over $15{,}000\,\rm{deg}^{2}$, primarily targeting $\rm{H\,\alpha}$] emitters in the redshift range $0.9<z<1.8$ [13]. Its Quick Data Release 1 is already publicly available [34]. The large cosmic volumes probed by DESI and Euclid enable sub-percent precision in cosmological measurements and require a rigorous understanding of systematic errors, particularly those related to redshift measurements.
In this work, the spectroscopic systematic redshift errors are referred to as redshift errors. The redshift errors exist in both ground-based and space-based galaxy redshift surveys. They include redshift uncertainty, the small statistical fluctuations in redshift measurements, and redshift catastrophic failures, the large errors from the true values. Both types of error persists in secure redshift measurements, i.e., those that have been calibrated and successfully passed the redshift determination pipeline. Consequently, these hidden errors can distort the observed clustering pattern and potentially bias our cosmological analyses. Since both redshift uncertainty and catastrophic failures can be quantified from repeated redshift observations, it is practical to analyze them together. This combined treatment is also relevant for space missions such as Euclid, which apply slitless spectroscopy to measure redshifts and are therefore more susceptible to both types of contamination. In the following, we will describe how redshift uncertainty and redshift catastrophic failures are characterized and modeled, and what are the redshift errors expected in slitless spectroscopic surveys.
@@ 2.1 Redshift uncertainty
Redshift uncertainty is an inevitable systematic error in spectroscopic redshift measurements. It arises because the spectral lines used to determine redshifts have specific widths rather than being ideal delta functions. These widths can result from a combination of physical effects (e.g., thermal motions, turbulence, gravitational broadening) and the instrumental resolution. Spectroscopic surveys report redshift uncertainty for different types of targets. For example, in eBOSS galaxy samples, the mean redshift uncertainty for LRG is $65.6\,\rm{km}\cdot s^{-1}$ [35]. For QSO, the typical uncertainty is around 300 $\rm{km}\cdot s^{-1}$ [36]. Repeat observations of ELG indicate that 50% of redshift estimates have a precision better than $20\,\rm{km}\cdot s^{-1}$ [37]. In the recent DESI Data Release 1 (DR1), the redshift precision reaches $\sim 10\,\rm{km}\cdot s^{-1}$ for BGS and ELG, $\sim 50\,\rm{km}\cdot s^{-1}$ for LRG, and $20$-$125\,\rm{km}\cdot s^{-1}$ for QSO in $0.8<z<1.8$ [27].
The effect of redshift uncertainty is equivalent to randomly smearing the peculiar velocity of an observed object along the line-of-sight. This smearing velocity is empirically quantified using redshift difference, $\Delta z$, from repeated observations, given by $\Delta v_{\rm{smear}}=c\,\Delta z/(1+\bar{z})$. The $c$ is the speed of light and $\bar{z}$ is the preferred redshift measurement of the object, typically taken as the mean or the highest-confidence one from repeats. The distribution of observed $\Delta v_{\rm{smear}}$ can be statistically modeled. For galaxies, this distribution is typically Gaussian $\mathcal{N}(\mu,\,\sigma_{\Delta v})$, with $\sigma_{\Delta v}$ characterizing the redshift uncertainty dispersion [35, 25, 26]:
@@
For QSO, the distribution of $\Delta v_{\rm{smear}}$ usually exhibits extended wide tails. In such case, a multiple-Gaussian model [36, 33] or a Lorentzian distribution [26], provides a better fit to repeat measurements. In our work, we use the Lorentzian profile, $\mathcal{L}(\,p,\,w_{\Delta v})$, to describe the QSO redshift uncertainty, which can be expressed as:
$\mathcal{L}(\Delta v_{\rm{smear}}\,|\,p,\,w_{\Delta v})=\frac{A}{1+((\Delta v_{\rm{smear}}-p)/w_{\Delta v})^{2}}$ (2.2)
Here, $w_{\Delta v}$ captures the width of the distribution, thus characterizing the redshift uncertainty in the Lorentzian profile. In our analyses, we adopt both Gaussian and Lorentzian profiles to model QSO redshift uncertainty. In this way, we can access how the presence of long-tailed uncertainty distributions may impact clustering and cosmological measurements.
@@ 2.2 Redshift catastrophic failures
Redshift catastrophic failures (or redshift catastrophics) arise in a small fraction of redshift measurements, typically identified through significant redshift differences in repeated observations. These errors can result from misidentification of sky residuals or confusion between spectral lines of different galaxies [14, 21, 17]. Faint or ambiguous spectral features can also lead to catastrophically incorrect redshift measurements. The catastrophics contamination rate, $f_{c}$, and the threshold for identifying it vary in different surveys. For example, the WiggleZ dark energy survey reports a 5% $f_{c}$ by defining redshift blunders with $|\Delta z|>400\,\rm{km}\cdot s^{-1}$ (larger than $4\sigma_{\Delta v}$) [14]. In eBOSS, the reported $f_{c}$ is less than 1% for LRGs and ELGs ($|\Delta z|>1000\,\rm{km}\cdot s^{-1}$) [37, 35] and 1.6% for quasars ($|\Delta z|>3000\,\rm{km}\cdot s^{-1}$) [36]. In DESI DR1, the $f_{c}$ is estimated to be less than 1% for all tracers at $z<2.1$ [27].
In the study of DESI DR1 repeated ELG observations, redshift catastrophics are classified into two main categories: sky confusion at $z=1.32$ and random catastrophics [21]. Sky confusion, responsible for 24% of these errors, arises from misidentifying residuals of sky emissions at around 8625 Å and can be approximated as a Gaussian distribution. The remaining 76% random catastrophics encompass various error patterns and are modeled using a log-normal profile as:
$\mathcal{LN}(\Delta\hat{v}_{\rm{catas}}\,|\,\mu_{\text{ran}},\sigma_{\text{ran}}^{2},v_{0})=\frac{A}{\sqrt{2\pi}\sigma_{\text{ran}}\Delta\hat{v}_{\rm{catas}}}\,\exp\left[-\frac{\left(\Delta\hat{v}_{\rm{catas}}-\mu_{\text{ran}}\right)^{2}}{2\sigma_{\text{ran}}^{2}}\right]$ (2.3)
This function models the absolute log-nomral distribution of redshift catastrophics ($\Delta\hat{v}_{\rm{catas}}=\log_{10}(|\Delta v_{\rm{catas}}|)$) with parameters $\mu_{\text{ran}}$, $\sigma_{\text{ran}}^{2}$ and $v_{0}$ to characterize. In our study, we focus on the random pattern while varying the contamination rate $f_{c}$ to generalize catastrophics patterns for different surveys.

### v4

@@ Abstract
Spectroscopic redshift errors, including redshift uncertainty and catastrophic failures, can bias cosmological measurements from galaxy clustering at sub-percent level. In this work, we investigate their impact on the full-shape analysis using contaminated mock catalogs. We find that redshift uncertainty introduces a scale-dependent damping effect on the power spectrum, which is absorbed by counterterms in clustering model, keeping parameter biases below $5\%$. Catastrophic failures suppress the power spectrum amplitude by an approximately constant factor, which scales with the catastrophic rate $f_{c}$. While this effect is negligible for DESI galaxy populations ($f_{c}=1\%$), the slitless-like errors, combining redshift uncertainty with $f_{c}=5\%$ catastrophics, introduce significant biases in cosmological constraints. In this case, we find $6\%$ to $16\%$ shifts ($\sim 2.2\sigma$ level) in estimating the fractional growth rate $df\equiv f/f^{\rm{fid}}$ and the log primordial amplitude $\ln(10^{10}A_{s})$. Applying the correction factor $(1-f_{c})^{2}$ on the galaxy power spectrum mitigates the bias but weakens the parameter constraints due to new degeneracies. Alternatively, fixing $f_{c}$ to its expected value restores the constraining power with a modest bias of $1.0\sigma$. Our results indicate that for space-based slitless surveys such as Euclid, at minimum accurate estimation of $f_{c}$ and its incorporation into the clustering model are necessary to get unbiased cosmological inference. Extending to evolving dark energy and massive neutrino cosmologies, redshift errors do not bias the dark energy properties parametrized by $w_{0}$ and $w_{a}$, but can degrade constraints on the summed neutrino mass $\sum m_{\nu}$ by up to 80% in the worst case.
@@ 1 Introduction
The large-scale structure (LSS) observed by galaxy redshift surveys has played an important role in advancing our understanding of fundamental physics. Clustering features, such as baryon acoustic oscillations (BAO; [eisenstein_baryonic_1998, eisenstein_detection_2005, weinberg_observational_2013, alam_clustering_2017, desi_collaboration_desi_2025-4]), and the full-shape power spectrum [ivanov_cosmological_2020, damico_cosmological_2020, desi_collaboration_desi_2025-1, desi_collaboration_desi_2025-3], have been used to constrain some key quantities in cosmology, including dark energy equation of state $w$, the matter density $\Omega_{\mathrm{m}}$, the structure growth rate $f\sigma_{8}$, neutrino mass $\sum m_{\nu}$ and possible deviations from general relativity. Current leading experiments, the Dark Energy Spectroscopic Instrument [^1] DESI, desi.lbl.gov (DESI; [desi_collaboration_desi_2016, desi_collaboration_desi_2016-1]), and the ongoing space mission Euclid [^2] Euclid, cosmos.esa.int/web/euclid [laureijs_euclid_2011, euclid_collaboration_euclid_2025-2] are mapping the three-dimensional Universe by measuring redshifts for tens to hundreds of millions galaxies. These surveys will cover over more than 15,000 $\rm deg^{2}$ of the sky, and trace the LSS back to redshift $\sim 3.5$. This large probing volume allows the clustering signal to be measured with minimal cosmic variance, leading to sub-percent precision in cosmological measurements. Given this level of precision, detailed studies on systematic errors are critical to ensure the robustness and reliability of cosmological constraints.
@@
Among various observational systematics, we focus on the impact of spectroscopic redshift errors. They persist in data with secure redshifts, where redshift calibration and algorithmic processing have already been applied [blake_wigglez_2010, bolton_spectral_2012, guy_spectroscopic_2023, euclid_collaboration_euclid_2025-1]. There are two primary types of errors: redshift uncertainty and redshift catastrophic failures. Redshift uncertainty refers to minor shifts ($\Delta z\sim 10^{-4}$) in all redshift measurements, primarily caused by instrumental noise and the intrinsic width of spectral features. In contrast, redshift catastrophic failures result from the misidentification of spectral lines, leading to severely incorrect redshifts ($\Delta z\sim 10^{-2}$) in a small fraction of data. These systematic errors can distort the observed clustering signal and potentially bias the cosmological inference. While previous studies have examined their individual impacts [blake_wigglez_2011, zarrouk_clustering_2018, hou_completed_2020, yu_elg_2025, krolewski_impact_2025, euclid_collaboration_euclid_2025-5], a comprehensive analysis that considers all these effects remains unexplored. This work aims to bridge that gap and explore their combined influence on cosmological inference from full-shape clustering analysis.
@@
The spectroscopic redshift errors can be quantified using repeated redshift observations and incorporated into realistic galaxy mocks (e.g., [smith_completed_2020, yu_model_2022, yu_desi_2023, yu_elg_2025]). By comparing analyses on clean mocks with those contaminated by redshift errors, we can assess their impact on clustering statistics and cosmological constraints. Improved facilities and increased statistics in DESI [desi_collaboration_data_2025, yu_elg_2025] provide a better way to model the redshift errors and study their impact on cosmological analyses. Additionally, space-based surveys such as Euclid use slitless spectroscopy at near-infrared to measure redshifts [euclid_collaboration_euclid_2025-4, euclid_collaboration_euclid_2025-3]. This method requires no pre-selection of targets and is particularly effective for probing high-redshift galaxies. However, it is also more susceptible to redshift errors due to the lower spectrum resolution and the potential overlap of different spectra. These challenges have been addressed in the context of redshift interlopers for Euclid [euclid_collaboration_euclid_2025-5]. In this work, we use hypothetical combined redshift errors to represent the conditions expected in slitless surveys. This approach provides a description of the related systematics and offers practical guidance for interpreting and mitigating their impact in cosmological analyzes for spectroscopic surveys.
@@
The outline of the paper is as follows: section˜2 introduces two galaxy redshift surveys, DESI and Euclid, and the spectroscopic redshift errors. section˜3 describes how we construct our contaminated mock catalogs. section˜4 presents the methodologies used to derive cosmological inferences from the mock data. In section˜5, we present the impact of redshift errors on clustering statistics and how these propagate into the cosmological parameter constraints, including extensions $w_{0}w_{a}\rm{CDM}$ and massive neutrino. Finally, we summarize our findings and conclude in section˜6.
@@ 2 Spectroscopic Survey and Redshift Errors
Spectroscopic surveys measure the redshifts of objects with high precision by analyzing their spectral features, such as emission and absorption lines. Two massive ongoing spectroscopic surveys, DESI and Euclid, are measuring galaxy clustering with unprecedented accuracy. In its first three years, the ground-based survey DESI has observed over 14 million galaxies [desi_collaboration_desi_2025-4] and plans to extend its initial plan to 60M galaxies across $18{,}000\,\rm{deg}^{2}$. Its dark matter tracers include Bright Galaxies (BGS) at $0.1<z<0.4$ [hahn_desi_2023], Luminous Red Galaxies (LRG) at $0.4<z<1.1$ [zhou_target_2023], Emission Line Galaxies (ELG, mainly [O ii] emitters) at $0.6<z<1.6$ [raichoor_target_2023], and Quasi-Stellar Objects (QSO) at $0.8<z<3.5$ [chaussidon_target_2023]. In parallel, the Euclid mission is conducting slitless spectroscopic survey from space and aim to observe tens of millions of galaxies over $15{,}000\,\rm{deg}^{2}$. Its primary targets are $\rm{H\,\alpha}$] emitters in the redshift range $0.9<z<1.8$ [euclid_collaboration_euclid_2025-2]. The Euclid Quick Data Release 1 (QR1) is already publicly available [euclid_collaboration_euclid_2025]. The large cosmic volumes probed by DESI and Euclid enable sub-percent precision in cosmological measurements and require a rigorous understanding of systematic errors, particularly those related to redshift measurements.
In this work, the spectroscopic systematic redshift errors are referred to as redshift errors. The redshift errors exist in both ground-based and space-based galaxy redshift surveys. These include redshift uncertainty, the small statistical fluctuations in redshift measurements, and redshift catastrophic failures, the large errors from the true values. Both types of error exists in secure redshift measurements, i.e., those that have been calibrated and successfully passed the redshift determination pipeline. Consequently, these hidden errors can distort the observed clustering pattern and potentially bias our cosmological analyses. Since both redshift uncertainty and catastrophic failures can be quantified from repeated redshift observations, e.g. $dv_{ij}=c(z_{i}-z_{j})/(1+z_{i})$, it is practical to analyze them together. This combined treatment is also relevant for space missions such as Euclid, which apply slitless spectroscopy to measure redshifts and are more susceptible to both types of contamination. In the following, we will describe how redshift uncertainty and redshift catastrophic failures are characterized and modeled, and what are the redshift errors expected in slitless spectroscopic surveys.
@@ 2.1 Redshift uncertainty
[…]
The effect of redshift uncertainty is equivalent to randomly smearing the peculiar velocity of an observed object along the line-of-sight. This smearing velocity, $\Delta v_{\rm{smear}}$, defined as the same as $dv_{ij}$, is empirically quantified using redshift difference. The distribution of observed $\Delta v_{\rm{smear}}$ can be statistically modeled. For galaxies, this distribution is typically Gaussian $\mathcal{N}(\mu,\,\sigma_{\Delta v})$, with $\sigma_{\Delta v}$ characterizing the redshift uncertainty dispersion [ross_completed_2020, yu_model_2022, yu_desi_2023]:
@@
For QSO, the distribution of $\Delta v_{\rm{smear}}$ usually exhibits extended wide tails. In such case, a multiple-Gaussian model [lyke_sloan_2020, chaussidon_target_2023] or a Lorentzian distribution [yu_desi_2023] provide a better fit to repeat observations. In our work, we use the Lorentzian profile, $\mathcal{L}(\,p,\,\mathrm{w}_{\Delta v})$, to describe the QSO redshift uncertainty with extended tail:
$\mathcal{L}(\Delta v_{\rm{smear}}\,|\,p,\,\mathrm{w}_{\Delta v})=\frac{1}{1+((\Delta v_{\rm{smear}}-p)/\mathrm{w}_{\Delta v})^{2}}$ (2.2)
Here, $\mathrm{w}_{\Delta v}$ captures the width of the distribution, thus characterizing the redshift uncertainty in the Lorentzian profile. To assess how the presence of long-tailed uncertainty distributions may impact clustering and cosmological measurements, we adopt both Gaussian and Lorentzian profiles to model QSO redshift uncertainty. A comparison between these profiles is shown in the left panel of Figure˜1.
@@ 2.2 Redshift catastrophic failures
Redshift catastrophic failures (or redshift catastrophics) arise in a small fraction of redshift measurements, typically identified through significant redshift differences in repeated observations. These errors can result from misidentification of sky residuals or confusion between spectral lines of different galaxies [blake_wigglez_2010, yu_elg_2025, euclid_collaboration_euclid_2025-1]. Faint or ambiguous spectral features can also lead to catastrophically incorrect redshifts. The catastrophic contamination rate, $f_{c}$, and the threshold for identifying it vary in different surveys. For example, the WiggleZ survey reports 5% $f_{c}$ of redshift blunders with $|\Delta z|>400\,\rm{km}\cdot s^{-1}$ (larger than $4\sigma_{\Delta v}$[blake_wigglez_2010]. In eBOSS, the reported $f_{c}$ is less than 1% for LRGs and ELGs ($|\Delta z|>1000\,\rm{km}\cdot s^{-1}$) and 1.6% for quasars ($|\Delta z|>3000\,\rm{km}\cdot s^{-1}$[raichoor_completed_2020, ross_completed_2020, lyke_sloan_2020]. For DESI DR1, using the same $|\Delta z|$ thresholds as in eBOSS, the $f_{c}$ is estimated to be less than 1% for all tracers at $z<2.1$ [desi_collaboration_data_2025].
In the study of DESI DR1 repeated ELG observations, redshift catastrophics are classified into two main categories: sky confusion at $z=1.32$ and random catastrophics [yu_elg_2025]. Sky confusion, responsible for 24% of these errors, arises from misidentifying residuals of sky emissions at around $8625\,\text{\AA }$ and can be approximated as a Gaussian distribution centered at the corresponding sky line positions. The remaining 76% are random catastrophic failures that encompass various error patterns. These can be modeled using a log-normal profile defined as follows:
$\mathcal{LN}(\Delta\hat{v}_{\rm{c}}|\,\mu_{\text{ran}},\sigma_{\text{ran}}^{2},\hat{v}_{0})=\frac{1}{\sqrt{2\pi}\sigma_{\text{ran}}(-\Delta\hat{v}_{\rm{c}}+\hat{v}_{0})}\,\exp\left\{-\frac{\left[\ln(-\Delta\hat{v}_{\rm{c}}+\hat{v}_{0})\!-\!\mu_{\text{ran}}\right]^{2}}{2\sigma_{\text{ran}}^{2}}\right\}$ (2.3)
This function models the absolute log-normal distribution of redshift catastrophics ($\Delta\hat{v}_{\rm{c}}=\log_{10}(|dv_{ij}|)$) with parameters $\mu_{\text{ran}}$, $\sigma_{\text{ran}}^{2}$ and $\hat{v}_{0}$. In our study, we focus on the random pattern while varying the contamination rate $f_{c}$ to generalize catastrophic patterns for different surveys. We note that including sky confusion would not affect significantly the results [yu_elg_2025]. The distributions of catastrophics for $f_{c}=1\%$ and $5\%$ are shown in the right panel of Figure˜1, with the magnitude of $\Delta v_{\rm{c}}$ varying in $10^{3}$$10^{6}\,\rm{km}\cdot s^{-1}$.